Find the inverse of the function:
g(x)=2x-3/4+x

Answer by CubeyThePenguin(3113)

a) y = (2x-3)/(4+x)
4y + xy = 2x - 3
xy - 2x = -4y - 3
x(y - 2) = -4y - 3
x = (-4y - 3)/(y - 2)

g^(-1)(x) = (-4x - 3)/(x - 2)